OpenStudy (anonymous):

Choose the correct graph of the given system of equations. y − 2x = −1 x + 3y = 4

OpenStudy (anonymous):

graph of two lines, one with a positive slope and one with a negative slope, that intersect at the point (negative 1, 1)

OpenStudy (zzr0ck3r):

\(y-2x=-1\) \(x+3y=4\) Multiply the second one by \(2\) and we get \(y-2x=-1\) \(2x+6y=8\) now add them together \(7y=7\) so that \(y=1\) and now we find \(x\). \(x+3(1)=4\implies x=1\)

OpenStudy (zzr0ck3r):

\((-1,1)\) is not their intersection. Make sense?

OpenStudy (anonymous):

give me a moment please

OpenStudy (zzr0ck3r):

no rush m8

OpenStudy (anonymous):

so its either the second or third one

OpenStudy (triciaal):

at the intersection the (x, y) given (-1, 1) satisfies both equations

OpenStudy (zzr0ck3r):

\(1-2(-1)\ne -1\)

OpenStudy (zzr0ck3r):

\((-1,1)\) Does NOT satisfy either equation @triciaal

OpenStudy (triciaal):

graph of two lines, one with a positive slope and one with a negative slope, that intersect at the point (negative 1, 1) I misunderstood then so is the question incomplete?

OpenStudy (anonymous):

Choose the correct graph of the given system of equations. y − 2x = −1 x + 3y = 4

OpenStudy (zzr0ck3r):

I think those are his answers :)

OpenStudy (anonymous):

is it correct

OpenStudy (triciaal):

@GiulianaRestrepo we can't see your choices. the solution is the graph where the (x, y) satisfies both equation

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

sketch each graph on the same grid and find the point where they meet

OpenStudy (triciaal):

the first equation y = 2x -1 hint slope 2 and point (0, -1) is on that line the 2nd eqn 3y = -x + 4 slope negative and when y = 0 x = 4 so (4, 0) is on that line can you sketch each line and find (x, y) where they meet?

OpenStudy (anonymous):

i already solved it thanks !

OpenStudy (triciaal):

ok close the question.